OpenStudy (cookiemonster18):

1.Solve | x-1/5|=3 A.x=-14/5 or x=16/5 B.x=14/5 or x=-16/5 C.x=-14/5 or x=-16/5 D.x=14/5 or x=16/5 ***** 2.The table shows the relationship between two variables. Which selection describes the relationship? X=1, 2, 3, 4 Y=2, -3, -8, -13 A.increasing; linear B.increasing; nonlinear C.decreasing; linear ***** D.decreasing; nonlinear please correct me if im wrong

OpenStudy (phantomcrow):

Is that \[\frac{ x-1 }{ 5 }=3\] ?

OpenStudy (cookiemonster18):

no @PhantomCrow I corrected it. (I edited it) please look up at #1 for the correct problem I had to fix it since I done it wrong :)

OpenStudy (phantomcrow):

So is it \[|x-\frac{ 1 }{ 5 }|=3\] ?

OpenStudy (phantomcrow):

If so, you can begin by negating the absolute value and adding a positive/negative sign to three.

OpenStudy (cookiemonster18):

yes :)

OpenStudy (phantomcrow):

We do this because the absolute value of anything has to be positive, regardless of the value within the bars. So we should account for both negative and positive values.

OpenStudy (phantomcrow):

After this, we split the equation up into two separate solutions and solve for x in each one.

OpenStudy (cookiemonster18):

how do I do that @PhantomCrow

OpenStudy (phantomcrow):

\[|x-\frac{ 1 }{ 5 }|=3\] \[x-\frac{ 1 }{ 5 }=\pm3\] \[x-\frac{ 1 }{ 5 }=3\] and \[x-\frac{ 1 }{ 5 }=-3\]

OpenStudy (cookiemonster18):

??? @PhantomCrow

OpenStudy (phantomcrow):

What don't you understand?

OpenStudy (phantomcrow):

Remember that the absolute value of anything must be positive. So no matter what we put into x with the absolute value bars around it, it ill output a positive number. There can only be two possible values for x-1/5 --3 and negative 3 because we are taking the absolute value of it and regardless of its sign, we will get a positive number so we should account for the negative values in the absolute value bars as well as the positive ones. So if the absolute value bars were not there, the answer could remain as 3 or potentially change to -3.

OpenStudy (cookiemonster18):

oh okay but for #1 I thought it was D and for #2 I thought it was C and im not sure if I was correct or not @PhantomCrow

OpenStudy (phantomcrow):

Did you solve the two equations we derived for the first one?

OpenStudy (crystal_bliss):

Break down the problem into these 2 equations \[x-1/5=3 -(x-1/5)=3\]

OpenStudy (crystal_bliss):

the first one is not D @CookieMonster18

OpenStudy (phantomcrow):

\[x=3+\frac{ 1 }{ 5 }\] \[x=-3+\frac{ 1 }{ 5 }\]

OpenStudy (mathmale):

| x-1/5|=3 can be broken down into two separate equations: \[x-\frac{ 1 }{ 5 }=3, and\]

OpenStudy (mathmale):

\[-(x-\frac{ 1 }{ 5 })=3\]

OpenStudy (cookiemonster18):

A?? because is x=16/5 -1/5 it'll equal 3

OpenStudy (phantomcrow):

We have two solutions for x.

OpenStudy (mathmale):

Eliminating the fraction 1/5 will make the solution easier. Multiply both equations by 5 to do this (multiply every term). Sove the resulting equations. You'll get two separate solutions.

OpenStudy (crystal_bliss):

yes it would be A you just forgot the negative

OpenStudy (mathmale):

Again, expect TWO solutions, and be sure to check both.

OpenStudy (cookiemonster18):

@Crystal_Bliss what do you mean I forgot the negative?? for #1 A?? you mean x=-14/5??

OpenStudy (cookiemonster18):

@pooja195 am I correct? #1 is A and #2 is C??